The formula for the perimeter of a square is P = 4s, where s is the length of each side. One square has a perimeter of 86 units, and its sides are 3 units longer than those of the second square. Which equation can be used to find x, the side of the second square? Please help, ive tried working out all the answers available and they all seem to come to a decimal.

Answer by ramkikk66(644)


The formula for the perimeter of a square is P = 4s, where s is the length of each side. One square has a perimeter of 86 units, and its sides are 3 units longer than those of the second square. Which equation can be used to find x, the side of the second square? Please  help, ive tried working out all the answers available and they all seem to come to a decimal.

Ans:
Perimeter of 1st square = 86
So, each side of 1st square = 86 / 4 = 21.5 units

Let x be the side of the 2nd square

x = 86/4 + 3 = 21.5 + 3 = 24.5 units

Yes, it is a decimal value because 86 is not a multiple of 4.

Hope you got it now :)